# Efficient Decoder for End-to-End Oriented Object Detection in Remote Sensing Images

### v1

@@ Abstract
Object instances in remote sensing images often distribute with multi-orientations, varying scales, and dense distribution. These issues bring challenges to end-to-end oriented object detectors including multi-scale features alignment and a large number of queries. To address these limitations, we propose an end-to-end oriented detector equipped with an efficient decoder, which incorporates two technologies, Rotated RoI attention (RRoI attention) and Selective Distinct Queries (SDQ). Specifically, RRoI attention effectively focuses on oriented regions of interest through a cross-attention mechanism and aligns multi-scale features. SDQ collects queries from intermediate decoder layers and then filters similar queries to obtain distinct queries. The proposed SDQ can facilitate the optimization of one-to-one label assignment, without introducing redundant initial queries or extra auxiliary branches. Extensive experiments on five datasets demonstrate the effectiveness of our method. Notably, our method achieves state-of-the-art performance on DIOR-R (67.31% mAP), DOTA-v1.5 (67.43% mAP), and DOTA-v2.0 (53.28% mAP) with the ResNet50 backbone.
Index Terms:
Oriented object detection, remote sensing images, end-to-end detectors, attention.
Figure 1: (a) Object instances in remote sensing images exhibit arbitrary orientation, different scales and dense distribution. (b) Visualization of RRoI attention without using CAM. RRoI attention align features and mostly attends to object extremities such as the front and back of ships, planes and vehicles.
I Introduction
@@
Oriented object detection aims to predict categories and a set of oriented boxes for each target and is widely applied in remote sensing images. Unlike horizontal object detection towards to natural scene, oriented object detection is still a challenging task due to the variety of oriented objects, e.g. arbitrary orientation, varying scales, and dense distribution. Moreover, modern oriented detection methods have achieved excellent results based on convolutional neural networks (CNNs). These methods can be divided into two types: single-stage [1, 2] and two-stage methods [3, 4]. Single-stage methods benefit from numerous anchors (position prior) and a simple architecture. Two-stage methods are trained by a coarse-to-fine style intrinsically to learn abundant features.
@@
Inspired by the attention and transformer in natural language processing (NLP) [5], many detectors introduce it into horizontal object detection to extract global features and show promising performance [6, 7]. Most end-to-end horizontal object detectors rely on an encoder-decoder framework and remove complicated hand-designed modules.
@@
However, when the end-to-end framework is integrated into oriented object detection, it will suffer from several challenges. First, objects appear in arbitrary orientations leading to misalignment between features and oriented boxes, shown in Fig. 1(a). Image features as values and oriented boxes as positional queries need to be aligned when they are interacted with in cross-attention. Second, the scales of objects in remote sensing is varying, shown in Fig. 1(a). Small and large objects require high and low resolution features respectively. Multi-scale features are necessary to adapt objects of different scales. Third, dense objects require a large number of queries. In remote sensing scenes [8], it’s common to find hundreds of densely packed object instances within a single image. In contrast, images in natural datasets [9] typically contain no more than one hundred objects, e.g. in the COCO dataset [9], nearly 99% of the images have fewer than 30 objects. So, 100 queries in DETR [6] are sufficient, but they are still fewer than the objects in some remote sensing images. To enhance performance, some horizontal detectors[10, 11] have increased the number of queries by more than 1000. However, if we apply the same rate of increase to remote sensing queries, computation overhead will be increased significantly. Last, similar queries hinder optimization. The end-to-end detectors adopt one-to-one label assignments. Consider the situation where two queries are entirely identical. When label assignment, one query is assigned as a positive sample, while the other is assigned as a background category. Such a problem can potentially introduce adverse effects into the training process [12].
@@
In this paper, we propose an end-to-end oriented object detector to address the above problems called Rotated RoI Transformer (RRoIFormer), which consists of two effective technologies, Rotated RoI attention (RRoI attention) and Selective Distinct Queries (SDQ). Specifically, the proposed RRoI attention can align image features with oriented objects through an attention mechanism. Learnable oriented boxes as positional queries extract aligned features which serve as the values in this attention mechanism. Content queries generate attention weights and aggregate the values naturally. The features around the center point of oriented boxes and within them are highlighted and scored by the attention mechanism, shown in Fig. 1(b). The time complexity of RRoI Attention is of a similar magnitude to deformable attention [7].
@@
In addition, our proposed Selective Distinct Queries (SDQ) collects queries from intermediate decoder layers once and then filters similar queries to obtain distinct queries. In modern end-to-end detectors, the decoder consists of 6 consecutive decoder layers [12, 13, 7] and each decoder layer predicts results supervised by loss functions. We have observed that queries in the intermediate layers can represent objects adequately compared to initial queries. Hence, we collect queries from intermediate layers as high-quality selective queries, without introducing a large number of initial queries [10] or extra auxiliary network branches [11, 14]. But there are a large number of similar queries [12] within the query set. These similar queries hinder the optimization of training. Our proposed method can significantly filter similar queries and keep the remaining queries distinct, which facilitates one-to-one label assignment and training. The number of distinct queries is dynamic and normally less than the initial ones. SDQ is embedded into intermediate decoder layers, rather than positioned before or after the entire decoder.
@@
    The Rotated RoI attention is proposed to align multi-scale image features and oriented boxes via a cross-attention mechanism for objects with arbitrary orientations and varying scales towards oriented object detection tasks.
  • $\bullet$
@@
    To avoid a large number of initial queries or extra branches, the proposed Selective Distinct Queries first collects sparse high-quality queries from the intermediate decoder layers. Second, it filters similar queries to obtain distinct queries, which can facilitate one-to-one label assignment and training.
  • $\bullet$
    Extensive experiments demonstrate the effectiveness of our approach. With the Resnet50 backbone, RRoIFormer achieves an mAP of 67.31% on DIOR-R, 67.43% on DOTA-v1.5, and 53.28% on DOTA-v2.0 respectively, establishing new state-of-the-art benchmarks.
II RELABED WORK
II-A Attention in Object Detection
@@
DETR [6] first introduces Transformer [5] into object detection, but the single feature map and vanilla attention are low efficiency to scale and large resolution images. The spatially modulated co-attention proposed in SMCA [15] combines vanilla attention with a Gaussian-like weight map to constrain co-attention to focus more on regions near the predicted box. The deformable attention modules proposed by Deformable DETR [7] only focus on some key sampling points around a reference. These key sampling points are responsible for both classification and regression, but at the supervision of orientation, they are located at special positions that are sub-optimal, like the catercorner and axis of boxes. Anchor DETR [16] designs Row-Column Decoupled Attention that performs the row attention and column attention successively. The decoupled row and column attention loses spatial and orientation information. Some works [17, 18] realize attention using dynamic convolution [19, 20]. Different from above attentions, our proposed Rotated RoI attention (RRoI attention) focus on interested regions of learnable boxes $B$ ($x,y,w,h,\theta$) distributed in multi-scale features and align the multi-scale features to oriented objects. The time complexity of RRoI Attention is the same order of magnitude as deformable attention.
II-B Methods of Query Update
@@
We classified current methods of query update into three types: the basic type, query enhancement type, and auxiliary decoder branch type. In all these types, queries are successively updated layer by layer in the decoder. In the basic way, a fixed number of queries which are initialized randomly are fed into the first decoder layer and the outputs are also the same number of initial queries. Subsequently, these queries are decoded successively by other decoder layers, each sharing the same structure [6, 13, 7]. Due to the sparsity and random initialization of the queries, some enhancement methods are proposed. Anchor DETR [16] attaches three patterns to each anchor point which enlarge query numbers. Dynamic Sparse R-CNN [21] replaces random initialization with a dynamic proposal generation mechanism to initialize proposals and features. DDQ [12] supervises dense boxes lying on features and selects top-scoring distinct proposals as inputs to the decoder. SQR [22] stacks queries from previous decoder layers and feeds them to later layers. DQ-Det [23] combines basic queries as modulated queries according to the convex combination. Some researchers introduce an auxiliary decoder branch to accelerate training convergence. To realize one-to-may label assignment, Group DETR [24] feeds multiple groups of queries to decoders as auxiliary branches in parallel, similar to the way in H-DETR [11]. DN-DETR [25] and DINO [10] add noises to bounding boxes and class labels as noising queries in an auxiliary branch to mitigate the instability of bipartite graph matching. Co-DETR [14] train multiple auxiliary heads supervised by one-to-many label assignments, such as ATSS [26] and Faster R-CNN [27]. Different from the above query learned way, our Selective Distinct Queries (SDQ) collects sparse high-quality queries from intermediate decoder layers which avoid a large number of initial queries or auxiliary decoder branches. To facilitate one-to-one label assignment, SDQ filters similar queries and obtains distinct queries. Moreover, SDQ is embedded among decoder layers, rather than served as an auxiliary head. Without losing generality, SDQ can be easily combined with other enhancement methods.
Figure 2: Rotated RoI attention. (a) Decoder with Rotated RoI attention. (b) Rotated RoI attention.
III PROPOSED METHOD
@@
In this paper, we propose an end-to-end oriented detector called RRoIFormer which is equipped with an efficient decoder. We first introduce the architecture in Sec. III-A. Specifically, we illustrate the Rotated RoI attention (RRoI attention) and Selective Distinct Queries (SDQ) in Sec. III-B and III-C respectively.
III-A Overall Architecture
Our RRoIFormer consists of a backbone and an efficient decoder shown in Fig. 2(a), where we use multi-scale feature maps to represent the backbone for brevity. Multi-scale features are extracted by backbone (e.g., ResNet50 [28], SwinTransformer [29]) and Feature Pyramid Network (FPN). Images in remote sensing scenes are very huge which leads to a large size of corresponding image features. To reduce the computational overhead, we do not introduce the encoder to further interact with features following [30, 31, 32]. The efficient decoder incorporates two technologies, Rotated RoI attention (RRoI attention) and Selective Distinct Queries (SDQ). The decoder consists of multiple successive decoder layers. Queries generated from every decoder layer are assigned by a one-to-one label assignment strategy and supervised by the same losses. Classification for categories and regression for positions of objects are two tasks in oriented object detection.
III-B Rotatted RoI Attention
@@
Decoder with Rotated RoI Attention. Both cross-attention and self-attention modules are in the decoder shown in Fig. 2(a). A small set of learnable content queries $Q\in\mathbb{R}^{N\times C}$ are high-dimension (e.g., $C=256$) latent vectors and are expected to encode semantic information of objects. For brevity, we use queries to represent content queries below. The positional encoding embeddings are randomly initialized with the same shape as queries and then added by queries following [6]. The queries and positional encodings are jointly trained with the networks. In self-attention modules, queries interact with each other and the outputs keep the same shapes as the inputs. In our proposed Rotated RoI attention, as cross-attention, queries interact with corresponding regions of aligned features. The inputs of RRoI attention include queries from self-attention, learnable oriented boxes as positional queries, and multi-scale feature maps. Though the queries represent object characteristics, we introduce learnable oriented boxes $B\in\mathbb{R}^{N\times 5}$ to localize the objects. These oriented boxes are represented by parameters ($x,y,w,h,\theta$), denoting the center coordinate, width, height, and angle, respectively. The number of oriented boxes always equals queries. Both queries and oriented boxes are initialized randomly as learnable embeddings [6, 13, 33] or come from enhancement modules [12, 7, 21, 16]. The multi-scale feature maps $\{x^{l}\}_{l=1}^{L}$, where $x^{l}\in\mathbb{R}^{C\times H_{l}\times W_{l}}$, are extracted by backbone (e.g., ResNet50 [28], SwinTransformer [29]) and Feature Pyramid Network (FPN). In the next section, we will discuss RRoI Attention modules specifically.
@@
Multi-Scale Rotated RoI Attention Module. Given oriented boxes $B\in\mathbb{R}^{N\times 5}$ and multi-scale feature maps $\{x^{l}\}_{l=1}^{L}$ in Fig. 2(b), to solve the misalignment problems, we first align the features as values $V$,
$V={\rm Concat}\{R_{l}(x^{l},\phi(B_{n}),r)\}_{n=1}^{N}$ (1)
@@
where integer $l$ index the input feature level, and $n$ index the oriented boxes. Function $\phi(B_{n})$ maps oriented boxes to different feature levels by their scales. Function $R_{l}(\cdot)$ warps region features from the corresponding feature $x^{l}$ precisely according to oriented boxes and pooling size $r$, e.g. Rotate RoI Align [34]. Next, the values $V\in\mathbb{R}^{N\times C\times r^{2}}$ will further interact with queries. Given queries $Q\in\mathbb{R}^{N\times C}$, the multi-scale Rotated RoI Attention is calculate by
@@
where $n$ indexes the oriented boxes and queries, $m$ indexes the attention heads and $k$ indexes all $r^{2}$ points at a single value map. The attention weight $A$ is obtained via linear projection over the queries. Specifically, the queries $Q$ are first fed to a linear projection operator of $M\times r^{2}$ channels and then are fed to a softmax operator to obtain the attention weight $A\in\mathbb{R}^{N\times M\times r^{2}}$. $A_{nmk}$ denotes the attention weight of the $k^{th}$ point at a single value map in the $m^{th}$ head. The scalar attention weight $A_{nmk}$ is normalized by $\sum_{k=1}^{r^{2}}A_{nmk}=1$. The $W_{m}^{\prime}$ and $W_{m}$ are input value projection and output projection matrix at $m^{th}$ head respectively.
Complexity For Rotated Rotated RoI Attention. When calculating the values, to align features for oriented boxes, the bilinear interpolation will be used. Four adjacent pixel values interpolate one precise target point. So the complexity for calculating values is $O(4Nr^{2}s)$, where $s$ denotes the sampling ratio and $r^{2}s$ denotes all sampling points in one oriented box. Then for projecting values to multiple heads, the complexity is $O(r^{2}NC^{2})$. Due to attention weights obtained via linear projection over queries, the complexity is $O(MNCr^{2})$. When attention weights and values interact, it is $O(NCr^{2})$. Finally, for aggregating sampled values, the complexity is $O(NC^{2})$. So the overall complexity of Rotated RoI attention is $O(4Nr^{2}s+r^{2}NC^{2}+NMCr^{2}+NCr^{2}+NC^{2})$. In our experiments, $M=8$, $r=7$, $s=4$ and $C=256$ by default, thus $r^{2}<C$ and the complexity is of $O(N(50C^{2}+441C+784))$.
@@
III-C Selective Distinct Queries
Basic Query Update in Decoder. Before the first decoder layer, the queries and oriented boxes are initialized as inputs to the decoder, shown in Fig. 3(a). For brevity, both initial queries and oriented boxes are denoted by $q^{0}=\{q_{1}^{0},q_{2}^{0},\cdot\cdot\cdot,q_{n}^{0}\}$ here. In the basic query update mode, there are $d$ identical decoder layers (e.g. $d=6$) stacked successively. Each decoder layer $D^{i}$ takes the queries $q^{i-1}$ from the previous layer and processes them to obtain the updated queries $q^{i}$ for the next layer,

### v4

@@ Abstract
Oriented object detection presents a challenging task due to the presence of object instances with multiple orientations, varying scales, and dense distributions. Recently, end-to-end detectors have made significant strides by employing attention mechanisms and refining a fixed number of queries through consecutive decoder layers. However, existing end-to-end oriented object detectors still face two primary challenges: 1) misalignment between positional queries and keys, leading to inconsistency between classification and localization; and 2) the presence of a large number of similar queries, which complicates one-to-one label assignments and optimization. To address these limitations, we propose an end-to-end oriented detector called the Rotated Query Transformer, which integrates two key technologies: Rotated RoI Attention (RRoI Attention) and Selective Distinct Queries (SDQ). First, RRoI Attention aligns positional queries and keys from oriented regions of interest through cross-attention. Second, SDQ collects queries from intermediate decoder layers and filters out similar ones to generate distinct queries, thereby facilitating the optimization of one-to-one label assignments. Finally, extensive experiments conducted on four remote sensing datasets and one scene text dataset demonstrate the effectiveness of our method. To further validate its generalization capability, we also extend our approach to horizontal object detection.
keywords:
[…]
1 Introduction
@@
Oriented object detection (Wen et al., 2023) aims to predict categories and a set of oriented bounding boxes for each target, and it is widely used in remote sensing (Chen et al., 2023c; Xiao et al., 2023; Zhou et al., 2022a) and scene text detection. Unlike horizontal object detection, which focuses on natural scenes, oriented object detection remains a challenging task due to the diversity of oriented objects, such as arbitrary orientations, varying scales, and dense distributions. Although modern oriented detection methods, particularly those based on convolutional neural networks (CNNs), have achieved excellent results, challenges persist. These methods can be categorized into two types: single-stage (Xu et al., 2023; Yu and Da, 2023) and two-stage (Xie et al., 2021; Han et al., 2021b) approaches. Single-stage methods benefit from numerous anchors (position priors) and a straightforward architecture, while two-stage methods are inherently designed to learn rich features through a coarse-to-fine training style.
@@
Inspired by the attention and transformer in natural language processing (NLP) (Vaswani et al., 2017), many detectors introduce it into horizontal object detection to extract global features and show promising performance (Carion et al., 2020; Zhu et al., 2020). Most end-to-end horizontal object detectors rely on an encoder-decoder framework and remove complicated hand-designed modules.
@@
However, when the end-to-end framework is integrated into oriented object detection, it will suffer from several challenges. First, objects appear in arbitrary orientations leading to misalignment between positional queries and keys, shown in Fig. 1(a). Image features as values and oriented boxes as positional queries need to be aligned when they are interacted with in cross-attention. In addition, the scales of objects vary. Small and large objects require high and low resolution features respectively. Therefore values are extracted from multiple levels of features to adequately represent objects. Second, similar queries hinder optimization. The end-to-end detectors adopt one-to-one label assignments. Consider the situation where two queries are entirely identical. When label assignment, one query is assigned as a positive sample, while the other is assigned as a background category. Such a problem can potentially introduce adverse effects into the training process (Zhang et al., 2023b). To alleviate it, some horizontal detectors (Zhang et al., 2023a; Jia et al., 2023) have increased the number of queries by more than 1000. However, it’s common to find hundreds of densely packed object instances within a single image in oriented object detection. Dense objects require a large number of queries. Applying the same rate of increase to dense object queries will result in a significant increase in computational overhead.
Figure 1: (a) Object instances in oriented object detection exhibit arbitrary orientation, different scales and dense distribution. (b) Visualization of RRoI attention without using CAM. RRoI attention align features and mostly attends to object extremities such as the front and back of ships, planes and vehicles.
@@
In this paper, we propose an end-to-end oriented object detector to address the above problems called Rotated Query Transformer (RQFormer), which consists of two effective technologies, Rotated RoI attention (RRoI attention) and Selective Distinct Queries (SDQ). Specifically, the proposed RRoI attention can align keys with positional queries through an attention mechanism. Learnable oriented boxes as positional queries extract aligned features that serve as the values in this attention mechanism. Content queries generate attention weights and aggregate the values naturally. The features around the center point of positional queries and within them are highlighted and scored by the attention mechanism, shown in Fig. 1(b). The time complexity of RRoI Attention is of a similar magnitude to deformable attention (Zhu et al., 2020).
@@
In addition, our proposed Selective Distinct Queries (SDQ) collects queries from intermediate decoder layers once and then filters similar queries to obtain distinct queries. In modern end-to-end detectors, the decoder consists of 6 consecutive decoder layers (Zhang et al., 2023b; Sun et al., 2021; Zhu et al., 2020) and each decoder layer predicts results supervised by loss functions. We have observed that queries in the intermediate layers can represent objects adequately compared to initial queries. Hence, we collect queries from intermediate layers as high-quality selective queries, without introducing a large number of initial queries (Zhang et al., 2023a) or extra auxiliary network branches (Jia et al., 2023; Zong et al., 2023). These queries can effectively cover objects. But there are still a large number of similar queries (Zhang et al., 2023b) within the query set. These similar queries hinder the optimization of training. Our method can significantly filter similar queries and keep the remaining queries distinct, which facilitates one-to-one label assignment and training. The number of distinct queries is dynamic and normally less than the initial ones. SDQ is embedded into intermediate decoder layers, rather than positioned before or after the entire decoder.
@@
  • To address the misalignment, the Rotated RoI attention is proposed to align positional queries and keys via a cross-attention mechanism for objects with arbitrary orientations and varying scales towards oriented object detection tasks.
@@
  • To mitigate the issue of query similarity, the proposed Selective Distinct Queries first collects sparse high-quality queries from the intermediate decoder layers. Second, it filters similar queries to obtain distinct queries, which can facilitate one-to-one label assignment and training.
  • Extensive experiments on six datasets demonstrate the effectiveness of our approach, including oriented object detection, text detection, and horizontal object detection.
2 Related Work
2.1 End-to-end Oriented Object Detection
End-to-end transformer-based methods (Carion et al., 2020; Zhu et al., 2020; Zhang et al., 2023a) have demonstrated promising performance in horizontal object detection tasks. Inspired by them, some researchers have extended Transformer-based approaches to oriented object detection. AO2-DETR (Dai et al., 2022) adapt Deformable DETR (Zhu et al., 2020) to oriented object detection task. It designs oriented box generation and refinement modules to provide accurate position prior. ARS-DETR (Zeng et al., 2024) specializes in adapting DN-DETR (Li et al., 2022a) for the oriented object detection task. It proposes the rotated deformable attention, where sampling points are rotated based on angles for feature alignment. RHINO (Lee et al., 2023) takes a route by adapting DINO (Zhang et al., 2023a) for oriented object detection and improves L1 distance cost to Hausdorff distance cost in Hungarian Matching of one-to-one label assignment for more stable training. In addition, several studies have focused on enhancing one-to-one label assignment. EMO2-DETR (Hu et al., 2023) observes and addresses the issue that one-to-one label assignment results in relative redundancy of object queries because objects are unevenly distributed in images. Furthermore, certain methods focus on improving object queries. PSD-SQ (Feng and Wang, 2024) represents object queries as point sets rather than oriented boxes for accurate instance feature sampling. D2Q-DETR (Zhou et al., 2023) designs dynamic queries which gradually reduce the number of object queries in the stacked decoder layers to better balance model precision and efficiency. Different from the above methods, our proposed approach introduces Rotated RoI attention for alignment between positional queries and keys and Selective Distinct Queries to facilitate one-to-one label assignment.
2.2 Attention in Object Detection
@@
DETR (Carion et al., 2020) first introduces Transformer (Vaswani et al., 2017) into object detection, but the single feature map and vanilla attention are low efficiency to scale and large resolution images. The spatially modulated co-attention proposed in SMCA (Gao et al., 2021) combines vanilla attention with a Gaussian-like weight map to constrain co-attention to focus more on regions near the predicted box. The deformable attention modules proposed by Deformable DETR (Zhu et al., 2020) only focus on some key sampling points around a reference. These key sampling points are responsible for both classification and regression, but at the supervision of orientation, they are located at special positions that are sub-optimal, like the catercorner and axis of boxes. Anchor DETR (Wang et al., 2022) designs Row-Column Decoupled Attention that performs the row attention and column attention successively. The decoupled row and column attention lose spatial and orientation information. Some works (Dai et al., 2021) realize attention using dynamic convolution. Different from the above attentions, our proposed Rotated RoI attention (RRoI attention) focuses on interested regions of positional queries $B$ ($cx,cy,w,h,\theta$) distributed in multi-scale features and aligns keys to oriented objects. The time complexity of RRoI Attention is the same order of magnitude as deformable attention.
Figure 2: Overall framework. Decoder incorporates Rotated RoI attention (RRoI attention) and Selective Distinct Queries (SDQ). RRoI attention serves as cross-attention, aligning keys with positional queries. SDQ collects queries from intermediate decoder layers and filters similar queries.
2.3 Methods of Query Update
@@
We classified current methods of query update into three types: the basic type, query enhancement type, and auxiliary decoder branch type. In all these types, queries are successively updated layer by layer in the decoder. In the basic way, a fixed number of queries which are initialized randomly are fed into the first decoder layer and the outputs are also the same number of initial queries. Subsequently, these queries are decoded successively by other decoder layers, each sharing the same structure (Carion et al., 2020; Sun et al., 2021; Zhu et al., 2020). Due to the sparsity and random initialization of the queries, some enhancement methods are proposed. Anchor DETR (Wang et al., 2022) attaches three patterns to each anchor point which enlarges query numbers. Dynamic Sparse R-CNN (Hong et al., 2022) replaces random initialization with a dynamic proposal generation mechanism to initialize proposals and features. DDQ (Zhang et al., 2023b) supervises dense boxes lying on features and selects top-scoring distinct proposals as inputs to the decoder. SQR (Chen et al., 2023a) stacks queries from previous decoder layers and feeds them to later layers. DQ-Det (Cui et al., 2023) combines basic queries as modulated queries according to the convex combination. Some researchers introduce an auxiliary decoder branch to accelerate training convergence. To realize one-to-may label assignment, Group DETR (Chen et al., 2023b) feeds multiple groups of queries to decoders as auxiliary branches in parallel, similar to the way in H-DETR (Jia et al., 2023). DN-DETR (Li et al., 2022a) and DINO (Zhang et al., 2023a) add noises to bounding boxes and class labels as noising queries in an auxiliary branch to mitigate the instability of bipartite graph matching. Co-DETR (Zong et al., 2023) train multiple auxiliary heads supervised by one-to-many label assignments. Different from the above query learned way, our Selective Distinct Queries (SDQ) collects sparse high-quality queries from intermediate decoder layers which avoid a large number of initial queries or auxiliary decoder branches. To mitigate the issue of query similarity, SDQ filters similar queries and obtains distinct queries. Moreover, SDQ is embedded among decoder layers, rather than served as an auxiliary head. Without losing generality, SDQ can be easily combined with other enhancement methods.
3 PROPOSED METHOD
@@
In this paper, we propose an end-to-end oriented detector called RQFormer which is equipped with the Rotated RoI attention (RRoI attention) and Selective Distinct Queries (SDQ). We first introduce the overall architecture in Sec. 3.1. Specifically, we illustrate the RRoI attention and SDQ in Sec. 3.2 and 3.3, respectively. Last, we introduce losses in Sec. 3.4.
Figure 3: (a) Decoder layer with Rotated RoI attention. (b) Rotated RoI attention.
3.1 Overall Architecture
Our RQFormer consists of a backbone, a Feature Pyramid Network (FPN), multiple decoder layers with Rotated RoI attention, the Selective Distinct Queries (SDQ) module, label assignment, and losses, shown in Fig. 2.
Backbone and FPN. The inputs of the backbone are images $x_{img}\in\mathbb{R}^{3\times H_{0}\times W_{0}}$, where 3, $H_{0}$, and $W_{0}$ denote color channels, heights, and widths of original images respectively. The output multi-scale features $\{x^{l}\}_{l=1}^{L}$ are extracted by backbone (e.g., ResNet50 (He et al., 2016)) and then FPN, where $x^{l}\in\mathbb{R}^{C\times H_{l}\times W_{l}}$, $L$ denotes levels of features, and $C$ represents channels of features. The scales of objects in oriented object detection vary. Small objects need high-resolution features, while large ones benefit from low-resolution features. Therefore, we use multi-scale features instead of a single feature level.
The vast size of images in remote sensing scenes results in high-resolution image features. To mitigate the computational burden, we do not introduce a dedicated image encoder for further interaction with features, in line with (Lv et al., 2023; Li et al., 2023). Only the FPN is used to feature fusion.
Content Queries and Positional Queries. Learnable positional and content queries serve as inputs to the decoder for representing object instances. The learnable positional queries $B\in\mathbb{R}^{N\times 5}$ denote coordinates of objects as $(cx,cy,w,h,\theta)$, signifying the center coordinate, width, height, and angle, respectively. Content queries $Q\in\mathbb{R}^{N\times C}$ are learnable high-dimension latent vectors, capturing semantic information of objects, where $C$ denotes dimensions of content queries, numerically the same as the channel of image features. Our method initializes positional and content queries through DDQ (Zhang et al., 2023b). They undergo sequential updates within the decoder, layer by layer. The number $N$ of content queries always equals that of positional queries. Learnable positional queries handle regression, while content queries manage classification.
Self-attention in Decoder. Self-attention enables the model to inhibit duplicate predictions (Carion et al., 2020) and facilitates interaction between content queries. The inputs of self-attention are content queries $Q$ and positional encoding $PE\in\mathbb{R}^{N\times C}$. The positional encoding comprises learnable embeddings (Carion et al., 2020), initialized with the same shape as content queries. The self-attention is calculated as follows:
$\begin{split}{\rm MHSA}(Q,K_{q},V_{q})=\\ {\rm softmax}((Q&+PE)(K_{q}+PE)^{\top}/\sqrt{d_{q}})V_{q}\end{split}$ (1)
where $K_{q}$ and $V_{q}$ are generated from content queries, sharing the same dimensions. The $d_{q}$ denotes dimensions. The outputs of self-attention keep the same shapes as the inputs.
Cross-attention and SDQ in Decoder. Our decoder incorporates two technologies, Rotated RoI attention (RRoI attention) and Selective Distinct Queries (SDQ). RRoI attention serves as cross-attention, aligning positional queries with keys through an attention mechanism. The decoder comprises consecutive layers, with SDQ embedded between two intermediate layers. SDQ can avoid a large number of initial queries and facilitate one-to-one label assignment and training. In the next part, we will provide a detailed elaborate on these components.
3.2 Rotated RoI Attention
@@
Decoder with Rotated RoI Attention. Both cross-attention and self-attention modules are in the decoder, shown in Fig. 3(a). A small set of learnable content queries $Q\in\mathbb{R}^{N\times C}$ are high-dimension (e.g., $C=256$) latent vectors and are expected to encode semantic information of objects. The positional encoding embeddings are randomly initialized with the same shape as queries and then added by queries following (Carion et al., 2020). The content queries and positional encodings are jointly trained with the networks. In self-attention modules, content queries interact with each other and the outputs keep the same shapes as the inputs. In our proposed Rotated RoI attention, as cross-attention, content queries interact with corresponding regions of aligned features. The inputs of RRoI attention include content queries from self-attention, positional queries, and multi-scale feature maps. Though the content queries represent object characteristics, we introduce positional queries $B\in\mathbb{R}^{N\times 5}$ to localize the objects. These positional queries are represented by parameters ($cx,cy,w,h,\theta$), denoting the center coordinate, width, height, and angle, respectively. The number of positional queries always equals content queries. Both content and positional queries are initialized through DDQ (Zhang et al., 2023b). The multi-scale feature maps $\{x^{l}\}_{l=1}^{L}$, where $x^{l}\in\mathbb{R}^{C\times H_{l}\times W_{l}}$, are extracted by backbone (e.g., ResNet50 (He et al., 2016)) and Feature Pyramid Network (FPN). In the next part, we will discuss RRoI attention modules specifically.
@@
Multi-Scale Rotated RoI Attention Module. Given positional queries $B\in\mathbb{R}^{N\times 5}$ and multi-scale feature maps $\{x^{l}\}_{l=1}^{L}$ in Fig. 3(b), to solve the misalignment problems, we first align the features as values $V$,
$V={\rm Concat}\{\{R_{l}(x^{l},\phi(B_{n}),r)\}_{n=1}^{N}\}_{l=1}^{L}$ (2)
@@
where integer $l$ index the input feature level, and $n$ index the positional queries. The values $V\in\mathbb{R}^{N\times C\times r^{2}}$ will further interact with content queries. Function $\phi(B_{n})$ maps positional queries to different feature levels by their scales. Function $R_{l}(\cdot)$ warps region features from the corresponding feature $x^{l}$ precisely according to positional queries and pooling size $r$, e.g., Rotate RoI Align (Ding et al., 2019). The function $R_{l}(\cdot)$ is calculated as follows:
$\begin{split}&R_{l}(x^{l},\phi(B_{n}),r)=\mathrm{Concat}\{\{\mathrm{Bi}(x^{l}[\omega(i,j)])\}_{i=1}^{r}\}_{j=1}^{r},\\ &\omega(i,j)=\begin{pmatrix}\cos\theta_{n}&-\sin\theta_{n}\\ \sin\theta_{n}&\cos\theta_{n}\end{pmatrix}\begin{pmatrix}(cx_{n}-\frac{w_{n}}{2}+\frac{iw_{n}}{r})\frac{W_{l}}{W_{0}}\\ (cy_{n}-\frac{h_{n}}{2}+\frac{jh_{n}}{r})\frac{H_{l}}{H_{0}}\end{pmatrix}\end{split}$ (3)
The focused values $V$ are calculated using bilinear operator $\mathrm{Bi}(\cdot)$ on the corresponding feature $x^{l}$. The indices of the focused regions are calculated by $\omega(i,j)$. The concatenation is used because the images and features lie on a 2D plane, taking into account both the horizontal and vertical axis.
The attention weight $A$ is obtained via linear projection over the queries. The learnable attention weights help RRoI attention focus on different parts of the values and heads.
$A=softmax(Linear(Q))$ (4)
Specifically, the content queries $Q$ are first fed to a linear projection operator of $M\times r^{2}$ channels and then are fed to a softmax operator to obtain the attention weight $A\in\mathbb{R}^{N\times M\times r^{2}}$. The $M$ denotes the number of attention heads.
Given content queries $Q\in\mathbb{R}^{N\times C}$, the multi-scale Rotated RoI Attention is calculate by
@@
where $n$ indexes the positional and content queries, $m$ indexes the attention heads and $k$ indexes all $r^{2}$ points at a single value map. $A_{nmk}$ denotes the attention weight of the $k^{th}$ point at a single value map in the $m^{th}$ head. The scalar attention weight $A_{nmk}$ is normalized by $\sum_{k=1}^{r^{2}}A_{nmk}=1$. The $W_{m}^{\prime}$ and $W_{m}$ are input value projection and output projection matrix at $m^{th}$ head respectively.
Complexity For Rotated Rotated RoI Attention. When calculating the values, to align features for positional queries, the bilinear interpolation will be used. Four adjacent pixel values interpolate one precise target point. So the complexity for calculating values is $O(4Nr^{2}s)$, where $s$ denotes the sampling ratio and $r^{2}s$ denotes all sampling points in one positional query. Then for projecting values to multiple heads, the complexity is $O(r^{2}NC^{2})$. Due to attention weights obtained via linear projection over queries, the complexity is $O(MNCr^{2})$. When attention weights and values interact, it is $O(NCr^{2})$. Finally, for aggregating sampled values, the complexity is $O(NC^{2})$. So the overall complexity of Rotated RoI attention is $O(4Nr^{2}s+r^{2}NC^{2}+NMCr^{2}+NCr^{2}+NC^{2})$. In our experiments, $M=8$, $r=7$, $s=4$ and $C=256$ by default, thus $r^{2}<C$ and the complexity is of $O(N(50C^{2}+441C+784))$. The complexity is of a similar magnitude to deformable attention.
@@
3.3 Selective Distinct Queries
Basic Query Update in Decoder. Before the first decoder layer, the content and positional queries are initialized as inputs to the decoder, shown in Fig. 4(a). For brevity, both initial content and positional queries are denoted by $q^{0}=\{q_{1}^{0},q_{2}^{0},\cdot\cdot\cdot,q_{n}^{0}\}$ here. In the basic query update mode, there are $d$ identical decoder layers (e.g., $d=6$) stacked successively. Each decoder layer $D^{i}$ takes the queries $q^{i-1}$ from the previous layer and processes them to obtain the updated queries $q^{i}$ for the next layer,